OpenStudy (anonymous):

passes through A(-3, 0) and B(-6, 5). What is the equation of the line that passes through the origin and is parallel to ?

OpenStudy (anonymous):

do you know to find the slop of a ligne?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

this the standard form equation of a ligne

OpenStudy (anonymous):

5x – 3y = 0 -x + 3y = 0 -5x – 3y = 0 3x + 5y = 0 -3x + 5y = 0

OpenStudy (anonymous):

those are my choices

OpenStudy (anonymous):

m= rise over run therefore m= (y2-y1)/ (x2-x1) where y2=5 and y1=0 x2=-6 and x1=-3

OpenStudy (anonymous):

i got 5/9

OpenStudy (anonymous):

what is